If the density of some hydrogen gass is 0.911 g/L at 214 K and 23.77 kPa, what would its density be at STP/
1
Expert's answer
2019-06-11T11:25:16-0400

Solution.

"\rho = \frac{p \times M}{R \times T}"

p1 = 23.77 kPa

T1 = 214

p2 = 101.325 kPa

T2 = 273 K



"\frac{\rho 1}{\rho 2} = \frac{p1 \times M \times R \times T2}{p2 \times M \times R \times T1}"

"\frac{\rho 1}{\rho 2} = \frac{p1 \times T2}{p2 \times T1}"

"\rho 2 = \frac{\rho 1 \times p2 \times T1}{p1 \times T2}"

"\rho 2 = \frac{0.911 \times 101.325*10^3 \times 214}{23.77*10^3 \times 273}"

"\rho 2 = 3.044"

Answer:

"\rho 2 = 3.044 g/L"